# DeepLag: Discovering Deep Lagrangian Dynamics for Intuitive Fluid Prediction

### v4

@@ 1 Introduction
In parallel to the Eulerian method, we notice another major approach for elucidating fluid dynamics, the Lagrangian method [13], also known as the particle tracking method. This method primarily focuses on tracing individual fluid particles by modeling the temporal evolution of their position and velocity. Unlike the Eulerian methods, which observe fluid flow at fixed spatial locations, the Lagrangian approach describes the fluid dynamics through the moving trajectory of individual fluid particles, offering a more natural and neat representation of fluid dynamics with inherent advantages in capturing intricate flow dynamics. Moreover, it allows a larger inference time step than Eulerian methods while complying with the Courant–Friedrichs–Lewy condition [6] that guarantees stability. In Figure 1, we can find that fluid dynamics is much more visually apparent in Lagrangian trajectories on the left compared to the density changes observed on static Eulerian grids on the right.
@@
  • Inspired by Lagrangian mechanics, we present EuLag Block, which can accurately track particle movements and interactively utilize the Eulerian features and dynamic information in fluid prediction, enabling a better dynamic modeling paradigm.
@@ 2.1 Eulerian and Lagrangian Methods
While traditional particle-based (or mixed-representation) solvers demonstrate superior accuracy and adaptability in inferring small-scale phenomena and dealing with nonlinear and irregular boundary conditions, they require computing the acceleration of each particle through physical equations, followed by sequential updates of their velocity and position [31]. This pointwise modeling approach often demands a significant number of points to fully characterize the dynamics of the entire field to meet accuracy requirements. Moreover, the irregular arrangement of particles and difficulty in parallelization result in higher computational costs and challenges with particle interpolation and gridding [12]. This renders particle-based solvers suboptimal compared to Eulerian solvers, particularly in high-dimensional spaces and large-scale simulations. However, DeepLag leverages the strengths of both solvers, eschewing equations and instead utilizing Eulerian information to assist particle tracking directly. This greatly alleviates the pressure on Lagrangian representation, requiring significantly fewer representative particles to aggregate the dynamics of the entire field.
@@ 2.2 Neural Fluid Prediction
Since traditional CFD methods often require hours or even days for precise simulations [41], deep models have been widely explored as efficient surrogate models that can provide near-instantaneous predictions. These neural fluid prediction models approximate the solutions of governing fluid equations differently and can be categorized into three mainstream paradigms as in Figure 2(a-c).
@@
This branch of methods in Figure 2(b) adopts deep models to learn the mapping from coordinates to solutions and formalize PDE constraints along with initial and boundary conditions as the loss function [48, 29, 46, 47]. Though this paradigm can precisely approximate the PDE solution, they usually require exact formalization for coefficients and conditions, limiting their generality and applicability to real-world fluids that are usually partially observed [35]. Plus, the Eulerian input disables them from handling Lagrangian descriptions.
@@
Recently, a new paradigm, illustrated in Figure 2(c), has emerged where deep models learn the neural operators between input and target functions, e.g., past observations to future fluid predictions. Various neural operators have significantly advanced fluid prediction following the DeepONet [22], which directly approximates mappings between equation parameters and solutions. For Eulerian grid data, models based on U-Net [32] and ResNet [15] structures have been proposed [30, 27, 17], as well as variants addressing issues like generalizing to unseen domains [44], irregular mesh [11], and uncertainty quantification [51]. Transformer-based models [43] enhance modeling capabilities and efficiency by exploiting techniques like Galerkin attention [3], incorporating ensemble information from the grid [14], applying low-rank decomposition to the attention mechanism [20], and leveraging spectral methods in the latent space [52]. Additionally, FNO [21] learns mappings in the frequency domain, and MP-PDE [2] utilizes the message-passing mechanism. For Lagrangian fluid particle data, some CNN-based methods [34, 42] model particle interactions through redesigned basic modules, while GNN-based methods [33, 25] update particle positions using the Encode-Process-Decode paradigm. Despite the progress made by these methods, they are limited to one description and do not seamlessly combine Eulerian and Lagrangian views.
@@ 3 DeepLag
Following the convention of neural fluid prediction [21], we formalize the fluid prediction problem as learning future fluid given past observation, as shown in Figure 2(d). Given a bounded open subset of $d$-dimensional Euclidean space $\mathcal{D}\subset\mathbb{R}^{d}$ and an Eulerian space $\mathcal{U}\subset\mathbb{R}^{o}$ with $o$ observed physical quantities, let $\mathbf{u}_{t}(\mathbf{x})\subset\mathcal{U}$ and $\mathbf{u}_{t+1}(\mathbf{x})\subset\mathcal{U}$ represent the Eulerian fluid field observation at two consecutive time steps on a finite coordinate set $\mathbf{x}\subset\mathcal{D}$, we aim at fitting the mapping $\Phi:\mathbf{u}_{t}(\mathbf{x})\rightarrow\mathbf{u}_{t+1}(\mathbf{x})$. Concretely, provided initial $P$ step observations ${U}_{P}=\{{\mathbf{u}}_{1},\ldots,{\mathbf{u}}_{P}\}$, the fluid prediction process can be written as the following autoregressive paradigm:
$\begin{split}{U}_{t}=\{{\mathbf{u}}_{t-P+1},\ldots,{\mathbf{u}}_{t}\}&\xrightarrow[]{\mathcal{F}_{\theta}}\tilde{\mathbf{u}}_{t+1},\\ \end{split}$ (3)
where $t\geq P$ and $\mathcal{F}_{\theta}$ represents the learned mapping between ${U}_{t}$ and the predicted field $\tilde{\mathbf{u}}_{t+1}$.
Inspired by Eq. (2), we present DeepLag as an Eulerian-Lagrangian Recurrent Network, which utilizes EuLag Block to learn Eulerian features and important Lagrangian dynamics interactively at various scales to address the complex spatiotemporal correlations in fluid prediction. Specifically, we capture the temporal evolving features at fixed points from the Eulerian perspective and the spatial dynamic information of essential particles from the Lagrangian perspective through their movement. By integrating Lagrangian pivotal dynamic information into Eulerian features, we fully model the spatiotemporal evolution of the fluid field over time and motion. Moreover, DeepLag can obtain critical trajectories within fluid dynamics with high computational efficiency by incorporating high-dimensional Eulerian space into lower-dimensional Lagrangian space.
Figure 2: Three types of neural fluid prediction models (a-c) and overview of DeepLag (d). The EuLag Block accumulates the previous dynamics at each time and scale to guide the Eulerian field update and then evolves the particle movement and dynamic conditioned on the updated field.
@@ 3.1 Overall Framework
It is widely acknowledged that fluids exhibit different motion characteristics across varying scales [3, 52]. In order to capture the intrinsic dynamics information at different scales, we track the trajectories of the key particles on $L$ scales separately and propose an Eulerian-Lagrangian Recurrent Network to realize the interaction between Eulerian and Lagrangian information. For clarity, we omit the scale index $l$ for primary quantities in this subsection, where $l\in\{1,2,\cdots,L\}$.
@@
To better capture the complex dynamics within the fluid field, we sample crucial points from the observations in the Eulerian perspective to determine the initial position for Lagrangian tracking in the dynamic sampling module. Explicitly, for the first predicting step $t=P$, given observations ${\mathbf{u}}_{t}=\{{\mathbf{u}}_{t}(\mathbf{x}_{k})|\mathbf{x}_{k}\in\mathcal{D}_{l},1\leq k\leq N_{l}\}\in\mathbb{R}^{N_{l}\times o}$ on all $N_{l}$ points in the observation domain $\mathcal{D}_{l}\subset\mathbb{R}^{d}$ of each scale, we extract their spatial dynamics by a convolutional network. We then calculate the probability matrix $\mathbf{S}\in\mathbb{R}^{N_{l}}$ via softmax along the spatial dimension:
@@
where $\operatorname{ConvNet}(\cdot)$ consists of a convolutional, activation, and fully-connected layer. We then sample $M_{l}$ Lagrangian tracking particles using the probability matrix at each scale:
@@
where ${\mathbf{p}_{t}}=\{\mathbf{p}_{t,i}^{l}\in\mathcal{D}\}_{i=1}^{M_{l}}\in\mathbb{R}^{M_{l}\times d}$ represents the set of sampled $M_{l}$ particles.
@@
For the $t$-th step, we adopt a learnable embedding layer with multiscale downsampling $\operatorname{Down}(\cdot)$ to encode past observations or predictions $U_{t}=\{{\mathbf{u}}_{t-P+1},\ldots,{\mathbf{u}}_{t}\}$ to obtain the Eulerian representations ${\mathbf{u}}_{t}^{l}\in\mathbb{R}^{N_{l}\times C_{l}}$ at each scale.
Next, we integrate EuLag Block to fuse the Lagrangian dynamics at each scale, thus directing the evolution of Eulerian features. This interaction also enables Eulerian features to guide the progression of Lagrangian dynamics. For the $l$-th scale, we track $M_{l}$ key particles over time, with $\mathbf{p}_{t}=\{\mathbf{p}_{t,i}^{l}\in\mathcal{D}_{l}\}_{i=1}^{M_{l}}$ representing their positions and $\mathbf{h}_{t}=\{\mathbf{h}_{t,i}^{l}\in\mathbb{R}^{C_{l}}\}_{i=1}^{M_{l}}$ denoting their learned particle dynamics. As shown in Figure 2(d), the positions and dynamics of particles at the $l$-th scale are learned autoregressively using the EuLag Block, which can be written as:
$\begin{split}\tilde{\mathbf{u}}_{t+1},\{\mathbf{p}_{t+1}\},\{\mathbf{h}_{t+1}\}=\operatorname{EuLag}({\mathbf{u}}_{t},\{\mathbf{p}_{t}\},\{\mathbf{h}_{t}\}),\\ \end{split}$ (6)
where the scale index $l$ is omitted for simplicity and clarity in notation. The EuLag Block optimally leverages the complementary strengths of Eulerian and Lagrangian representations, facilitating enhanced training and mutual refinement between these two perspectives. More details about the specific implementation of the EuLag Block are expanded upon in the following subsection 3.2.
After getting the output Eulerian features $\tilde{\mathbf{u}}_{t+1}$ with Lagrangian particle position $\mathbf{p}_{t+1}$ and dynamics $\mathbf{h}_{t+1}$ at the $l$-th scale, we further aggregate $\tilde{\mathbf{u}}_{t+1}$ with the predicted Eulerian field at coarser scale by upsampling $\operatorname{Up}(\cdot)$. Eventually, the prediction $\tilde{\mathbf{u}}_{t+1}$ at step $t+1$ is decoded from $\tilde{\mathbf{u}}_{t+1}^{1}$ with a projection layer. We unfold the implementation of the overall architecture in Appendix A.2.
@@ 3.2 EuLag Block
As formalized in Eq. (6), we adopt a recurrent framework to interactively exploit the information from two fluid descriptions, which consist of two main components: Lagrangian-guided feature evolving and Eulerian-conditioned particle tracking. The following quantities are all in the $l$-th scale.
@@
Classical theories [9] and numerical algorithms [31] showed that fluid predictions can be solved by identifying the origin of the fluid parcel and interpolating the dependent variable from nearby grid points. However, without specifying certain PDEs, we cannot explicitly determine the former position of the particle on each Eulerian observed point. Thus, we first adaptively synthesize the Lagrangian dynamics of the tracked particles to guide the evolution of Eulerian features using a cross-attention mechanism. Expressly, we adopt a Lagrangian-to-Eulerian attention, where the Eulerian field ${\mathbf{u}}_{t}$ serves as queries, and Lagrangian dynamics with particle position $\mathbf{h}_{t}^{\prime}=\operatorname{Concat}(\mathbf{h}_{t},\mathbf{p}_{t})\in\mathbb{R}^{M_{l}\times(C_{l}+d)}$ is used as keys and values:
$\begin{split}\displaystyle\operatorname{LagToEu-Attention}({\mathbf{u}}_{t},{\mathbf{h}}_{t}^{\prime},{\mathbf{h}}_{t}^{\prime})=\operatorname{Softmax}\Big{(}\frac{\mathbf{W}_{Q}{\mathbf{u}}_{t}(\mathbf{W}_{K}\mathbf{h}^{\prime}_{t})^{T}}{\sqrt{C_{l}+d}}\Big{)}\mathbf{W}_{V}\mathbf{h}_{t}^{\prime},\end{split}$ (7)
where $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$ and $\mathbf{W}_{V}$ stand for linear projection layers. We wrap the above attention into a standard transformer block with residual as Lagrangian-dynamics-guided Eulerian Evolution process:
$\begin{split}\displaystyle\tilde{\mathbf{u}}_{t+1}={\mathbf{u}}_{t}+\operatorname{LagToEu}({\mathbf{u}}_{t},{\mathbf{h}}_{t}^{\prime},{\mathbf{h}}_{t}^{\prime}).\end{split}$ (8)
@@
Traditional Lagrangian methods rely on interactions among vast quantities of particles to estimate the future fluid fields. However, the high computational cost hinders the application in rapid, deep models. In other data-driven approaches, the sparse sampling of tracking particles is computational-friendly but cannot directly derive Lagrangian dynamics for every particle. Considering the equivalence of the Eulerian and Lagrangian representations indicated in Eq. (1), we propose to learn particle movements based on the Eulerian conditions. Concretely, we utilize another Eulerian-to-Lagrangian cross-attention, where the evolved Eulerian features of the whole fluid field are used to navigate the Lagrangian dynamics of sparse particles:
$\begin{split}\mathbf{p}_{t+1}&=\operatorname{TrmBlock}_{Q,K,V_{p}}(\mathbf{h}_{t}^{\prime},\tilde{\mathbf{u}}_{t+1})\\ \mathbf{h}_{\text{global},t+1}&=\operatorname{TrmBlock}_{Q,K,V_{h}}(\mathbf{h}_{t}^{\prime},\tilde{\mathbf{u}}_{t+1}),\end{split}$ (9)
where $\operatorname{TrmBlock}_{Q,K,V}(\cdot,\cdot)$ represents standard transformer block with $\mathbf{W}^{{}^{\prime}}_{Q},\mathbf{W}^{{}^{\prime}}_{K},\mathbf{W}^{{}^{\prime}}_{V}$ as linear projection matrix. The forecasted global Lagrangian dynamic and movement of tracking particles are produced by two transformers with the same query and key projection but different value projections. To better model the dynamic evolution of the particles, we gather local Lagrangian dynamic by employing bilinear interpolation to output Eulerian features $\tilde{\mathbf{u}}_{t+1}$ on new particle position $\mathbf{p}_{t+1}$, then use a linear function to interact it with global dynamics information ${\mathbf{h}}_{\text{global},t+1}$, naming ${\mathbf{h}}_{t+1}=\operatorname{Aggregate}(\tilde{\mathbf{u}}_{t+1},\mathbf{p}_{t+1},{\mathbf{h}}_{\text{global},t+1})$. Additionally, as the input flow field may have an open boundary, the particles could move out of the observation domain. We check the updated position of the particles and resample the ones that exit from the latest probability matrix $\mathbf{S}$. Similar to the previous Eulerian Evolution process, the above Lagrangian Tracking process can be summarized as:
$\begin{split}\displaystyle\mathbf{h}_{t+1},\mathbf{p}_{t+1}=\operatorname{EuToLag}({\mathbf{h}}_{t}^{\prime},\tilde{\mathbf{u}}_{t+1},\tilde{\mathbf{u}}_{t+1}).\end{split}$ (10)
Overall, the EuLag Block can fully utilize the complementary advantages of Eulerian and Lagrangian perspectives in describing fluid dynamics, thereby being better suited to fluid prediction. For more implementation details of the EuLag Block, please refer to Appendix A.3.

### v5

@@ 1 Introduction
In parallel to the Eulerian method, we notice another major approach for elucidating fluid dynamics, the Lagrangian method [13], also known as the particle tracking method. This method primarily focuses on tracing individual fluid particles by modeling the temporal evolution of their position and velocity. Unlike the Eulerian methods, which observe fluid flow at fixed spatial locations, the Lagrangian approach describes the fluid dynamics through the moving trajectory of individual fluid particles, offering a more natural and neat representation of fluid dynamics with inherent advantages in capturing intricate flow dynamics. Moreover, it allows a larger inference time step tha Eulerian methods while complying with the Courant–Friedrichs–Lewy condition [6] that guarantees stability. In Figure 1, we can find that fluid dynamics is much more visually apparent in Lagrangian trajectories on the left compared to the density changes observed on static Eulerian grids on the right.
@@
  • Inspired by Lagrangian mechanics, we present EuLag Block, which can accurately track particle movements and interactively utilize the Eulerian features and dynamic information in fluid prediction, enabling a better dynamics modeling paradigm.
@@ 2.1 Eulerian and Lagrangian Methods
While traditional particle-based (or mixed-representation) solvers demonstrate superior accuracy and adaptability in inferring small-scale phenomena and dealing with nonlinear and irregular boundary conditions, they require computing the acceleration of each particle through physical equations, followed by sequential updates of their velocity and position [31]. This pointwise modeling approach often demands a significant number of points to fully characterize the dynamics of the entire field to meet accuracy requirements. Moreover, the irregular arrangement of particles and difficulty in parallelization result in higher computational costs and challenges with particle interpolation and gridding [12]. This renders particle-based solvers suboptimal compared to Eulerian solvers, particularly in high-dimensional spaces and large-scale simulations. However, the proposed DeepLag leverages the strengths of both solvers, eschewing equations and instead utilizing Eulerian information to assist particle tracking directly. This greatly alleviates the pressure on Lagrangian representation, requiring significantly fewer representative particles to aggregate the dynamics of the entire field.
@@ 2.2 Neural Fluid Prediction
As computational fluid dynamics (CFD) methods often require hours or even days for simulations [41], deep models have been explored as efficient surrogate models that can provide near-instantaneous predictions. These neural fluid prediction models approximate the solutions of governing fluid equations differently and can be categorized into three mainstream paradigms as in Figure 2(a-c).
@@
This branch of methods in Figure 2(b) adopts deep models to learn the mapping from coordinates to solutions and formalize PDE constraints along with initial and boundary conditions as the loss functions [48, 29, 46, 47]. Though this paradigm can explicitly approximate the PDE solution, they usually require exact formalization for coefficients and conditions, limiting their generality and applicability to real-world fluids that are usually partially observed [35]. Plus, the Eulerian input disables them from handling Lagrangian descriptions.
@@
Recently, a new paradigm, illustrated in Figure 2(c), has emerged where deep models learn the neural operators between input and target functions, e.g., past observations to future fluid predictions. Since DeepONet [22], various neural operators have significantly advanced fluid prediction, which directly approximates mappings between equation parameters and solutions. For Eulerian grid data, models based on U-Net [32] and ResNet [15] architectures have been proposed [30, 27, 17], as well as variants addressing issues like generalizing to unseen domains [44], irregular mesh [11], and uncertainty quantification [51]. Transformer-based models [43] enhance modeling capabilities and efficiency by exploiting techniques like Galerkin attention [3], incorporating ensemble information from the grid [14], applying low-rank decomposition to the attention mechanism [20], and leveraging spectral methods in the latent space [52]. Additionally, FNO [21] learns mappings in the frequency domain, and MP-PDE [2] utilizes the message-passing mechanism. For Lagrangian fluid particle data, some CNN-based methods [34, 42] model particle interactions through redesigned basic modules, while GNN-based methods [33, 25] update particle positions using the Encode-Process-Decode paradigm. Despite the progress made by these methods, they are limited to one description and do not seamlessly combine Eulerian and Lagrangian views.
@@ 3 DeepLag
Following the convention of neural fluid prediction [21], we formalize the fluid prediction problem as learning future fluid given past observation, as shown in Figure 2(d). Given a bounded open subset of $d$-dimensional Euclidean space $\mathcal{D}\subset\mathbb{R}^{d}$ and a Eulerian space $\mathcal{U}\subset\mathbb{R}^{o}$ with $o$ observed physical quantities, letting $\mathbf{u}_{t}(\mathbf{x})\subset\mathcal{U}$ and $\mathbf{u}_{t+1}(\mathbf{x})\subset\mathcal{U}$ represent the Eulerian fluid field observation at two consecutive time steps on a finite coordinate set $\mathbf{x}\subset\mathcal{D}$, we aim at fitting the mapping $\Phi:\mathbf{u}_{t}(\mathbf{x})\rightarrow\mathbf{u}_{t+1}(\mathbf{x})$. Concretely, provided initial $P$ step observations ${U}_{P}=\{{\mathbf{u}}_{1},\ldots,{\mathbf{u}}_{P}\}$, the fluid prediction process can be written as the following autoregressive paradigm:
$\begin{split}{U}_{t}=\{{\mathbf{u}}_{t-P+1},\ldots,{\mathbf{u}}_{t}\}&\xrightarrow[]{\mathcal{F}_{\theta}}\mathbf{u}_{t+1},\\ \end{split}$ (3)
where $t\geq P$ and $\mathcal{F}_{\theta}$ represents the learned mapping between ${U}_{t}$ and the predicted field $\mathbf{u}_{t+1}$.
Inspired by the material derivative in Eq. (2), we present DeepLag as a Eulerian-Lagrangian Recurrent Network, which utilizes the EuLag Block to learn Eulerian features and Lagrangian dynamics interactively at various scales to address the complex spatiotemporal correlations in fluid prediction. Specifically, we capture the temporal evolving features at fixed points from the Eulerian perspective and the spatial dynamic information of essential particles from the Lagrangian perspective through their movement. By integrating Lagrangian pivotal dynamic information into Eulerian features, we fully model the spatiotemporal evolution of the fluid field over time and motion. Moreover, DeepLag can obtain critical trajectories within fluid dynamics with high computational efficiency by incorporating high-dimensional Eulerian space into lower-dimensional Lagrangian space.
Figure 2: Three types of neural fluid prediction models (a-c) and overview of DeepLag (d). The EuLag Block accumulates the previous dynamics at each time and scale to guide the Eulerian field update and then evolves the particle movement and dynamics conditioned on the updated field.
@@ 3.1 Overall Framework
It is widely acknowledged that fluids exhibit different motion characteristics across varying scales [3, 52]. In order to capture the intrinsic dynamics information at different scales, we track the trajectories of the key particles on $L$ scales separately and propose a Eulerian-Lagrangian Recurrent Network to realize the interaction between Eulerian and Lagrangian information. For clarity, we omit the scale index $l$ for primary physical quantities in this subsection, where $l\in\{1,2,\cdots,L\}$.
@@
To better capture complex dynamics in the fluid field, we sample by importance of the points from Eulerian perspective observations to determine initial positions for Lagrangian tracking. This is done by our dynamics sampling module. For the first predicting step $t=P$, given observations ${\mathbf{u}}_{t}=\{{\mathbf{u}}_{t}(\mathbf{x}_{k})|\mathbf{x}_{k}\in\mathcal{D}_{l},1\leq k\leq N_{l}\}\in\mathbb{R}^{N_{l}\times o}$ on all $N_{l}$ points in observation domain $\mathcal{D}_{l}\subset\mathbb{R}^{d}$ of each scale, we extract their spatial dynamics by a convolutional network. We then calculate the probability matrix $\mathbf{S}\in\mathbb{R}^{N_{l}}$ via softmax along spatial dimension:
@@
where $\operatorname{ConvNet}(\cdot)$ consists of a convolutional layer and a linear layer, with activation function in-between. We then sample $M_{l}$ Lagrangian tracking particles by the probability matrix at each scale:
@@
where ${\mathbf{p}_{t}}=\{\mathbf{p}_{t,i}^{l}\in\mathcal{D}_{l}\}_{i=1}^{M_{l}}\in\mathbb{R}^{M_{l}\times d}$ represents the set of sampled $M_{l}$ particles.
@@
For the $t$-th timestep, we adopt a learnable embedding layer with multiscale downsampling $\operatorname{Down}(\cdot)$ to encode past observations or predictions $U_{t}=\{{\mathbf{u}}_{t-P+1},\ldots,{\mathbf{u}}_{t}\}$ to obtain the Eulerian representations ${\mathbf{u}}_{t}^{l}\in\mathbb{R}^{N_{l}\times C_{l}}$ at each scale.
We integrate the EuLag Block to fuse the Lagrangian dynamics at each scale and direct the evolution of Eulerian features. This interaction simultaneously enables Eulerian features to guide the progression of Lagrangian dynamics. For the $l$-th scale, we track $M_{l}$ key particles over time, with $\mathbf{p}_{t}=\{\mathbf{p}_{t,i}^{l}\in\mathcal{D}_{l}\}_{i=1}^{M_{l}}$ representing their positions and $\mathbf{h}_{t}=\{\mathbf{h}_{t,i}^{l}\in\mathbb{R}^{C_{l}}\}_{i=1}^{M_{l}}$ denoting their learned particle dynamics. As shown in Figure 2(d), the positions and dynamics of the $M_{l}$ particles at the $l$-th scale are learned autoregressively using the EuLag Block, which can be written as
$\begin{split}\mathbf{u}_{t+1},\{\mathbf{p}_{t+1}\},\{\mathbf{h}_{t+1}\}=\operatorname{EuLag}({\mathbf{u}}_{t},\{\mathbf{p}_{t}\},\{\mathbf{h}_{t}\}),\\ \end{split}$ (6)
where the scale index $l$ is omitted for notation simplicity. The EuLag Block learns to optimally leverage the complementary strengths of Eulerian and Lagrangian representations, facilitating mutual refinement between these two perspectives towards an exceeding performance. More details about the specific implementation of the EuLag Block are elaborated in the following subsection 3.2.
After evolving into new Eulerian features $\mathbf{u}_{t+1}$, Lagrangian particle position $\mathbf{p}_{t+1}$ and dynamics $\mathbf{h}_{t+1}$ at the $l$-th scale, we further aggregate $\mathbf{u}_{t+1}$ with the predicted Eulerian field at a coarser scale by upsampling $\operatorname{Up}(\cdot)$. Eventually, the full-resolution prediction $\mathbf{u}_{t+1}$ at step $t+1$ is decoded from $\mathbf{u}_{t+1}^{1}$ with a projection layer. We unfold the implementation of the overall architecture in Appendix A.2.
@@ 3.2 EuLag Block
As stated in Eq. (6), we adopt a recurrent network to interactively exploit the information from two fluid descriptions, which consist of two main components: Lagrangian-guided feature evolving and Eulerian-conditioned particle tracking. The following quantities are all in the $l$-th scale.
@@
Classical theories [9] and numerical algorithms [31] show that fluid predictions can be solved by identifying the origin of the fluid parcel and interpolating the dependent variable from nearby grid points. However, without specifying certain PDEs, we cannot explicitly determine the former position of the particle on each Eulerian observed point. Thus, we first adaptively synthesize the Lagrangian dynamics of the tracked particles to guide the evolution of Eulerian features using a cross-attention mechanism. Formally, we adopt a Lagrangian-to-Eulerian cross-attention, where the Eulerian field ${\mathbf{u}}_{t}$ serves as queries, and Lagrangian dynamics concatenated with particle position $\mathbf{h}_{t}||\mathbf{p}_{t}\in\mathbb{R}^{M_{l}\times(C_{l}+d)}$ is used as keys and values:
$\begin{split}\displaystyle\operatorname{LagToEu-Attn}({\mathbf{u}}_{t},\mathbf{h}_{t}||\mathbf{p}_{t},\mathbf{h}_{t}||\mathbf{p}_{t})=\operatorname{Softmax}\left(\frac{\mathbf{W}_{Q}{\mathbf{u}}_{t}(\mathbf{W}_{K}\cdot\mathbf{h}_{t}||\mathbf{p}_{t})^{\sf T}}{\sqrt{C_{l}+d}}\right)\mathbf{W}_{V}\cdot\mathbf{h}_{t}||\mathbf{p}_{t},\end{split}$ (7)
where $\mathbf{W}_{Q}$, $\mathbf{W}_{K}$ and $\mathbf{W}_{V}$ stand for linear projections. We wrap the attention in a Transformer block with residual connections to implement the Lagrangian-dynamics-guided Eulerian evolution process:
$\begin{split}\displaystyle\mathbf{u}_{t+1}={\mathbf{u}}_{t}+\operatorname{LagToEu}({\mathbf{u}}_{t},\mathbf{h}_{t}||\mathbf{p}_{t},\mathbf{h}_{t}||\mathbf{p}_{t}).\end{split}$ (8)
@@
Traditional Lagrangian methods rely on interactions among vast quantities of particles to estimate the future fluid fields. However, the high computational cost hinders the application in deep surrogate models. In other data-driven approaches, the sparse sampling of particles is computational-friendly but cannot directly derive Lagrangian dynamics for other particles. Considering the equivalence of the Eulerian and Lagrangian representations indicated by the material derivative in Eq. (1), we propose to learn particle movements based on the Eulerian conditions. Concretely, we utilize another Eulerian-to-Lagrangian cross-attention, where the evolved dense Eulerian features are used to navigate the Lagrangian dynamics of sparse particles:
$\begin{split}\displaystyle\operatorname{EuToLag-Attn}(\mathbf{h}_{t}||\mathbf{p}_{t},\mathbf{u}_{t+1},\mathbf{u}_{t+1})=\operatorname{Softmax}\bigg{(}\frac{(\mathbf{W}^{\prime}_{Q}\cdot\mathbf{h}_{t}||\mathbf{p}_{t})(\mathbf{W}^{\prime}_{K}\mathbf{u}_{t+1})^{\sf T}}{\sqrt{C_{l}}}\bigg{)}\mathbf{W}^{\prime}_{V}\mathbf{u}_{t+1},\end{split}$ (9)
where we use a different set of $\mathbf{W}^{\prime}_{Q}$, $\mathbf{W}^{\prime}_{K}$, and $\mathbf{W}^{\prime}_{V}$. Similarly, the Transformer block $\operatorname{EuToLag}$ wrapping this attention produces the change of forecasted global Lagrangian dynamics $\delta\mathbf{h}_{\text{global},t}$ and movement of tracking particles $\delta\mathbf{p}_{t}$, which leads to the next step by residual connections:
$\begin{split}\displaystyle\mathbf{h}_{\text{global},t+1}||\mathbf{p}_{t+1}=\mathbf{h}_{t}||\mathbf{p}_{t}+\operatorname{EuToLag}(\mathbf{h}_{t}||\mathbf{p}_{t},\mathbf{u}_{t+1},\mathbf{u}_{t+1}).\end{split}$ (10)
To better model the dynamic evolution of the particles, we gather local Lagrangian dynamic by employing bilinear interpolation to evolved Eulerian features $\mathbf{u}_{t+1}$ on new particle position $\mathbf{p}_{t+1}$, then use a linear function to aggregate it with global dynamics information ${\mathbf{h}}_{\text{global},t+1}$:
${\mathbf{h}}_{t+1}=\operatorname{Aggregate}(\operatorname{Interpolate}(\mathbf{u}_{t+1},\mathbf{p}_{t+1}),{\mathbf{h}}_{\text{global},t+1}).$ (11)
Additionally, particles could move out of the observation domain as the input field may have an open boundary. We check the updated position of tracking particles and resample from the latest probability matrix $\mathbf{S}$ to substitute the ones that exit, ensuring the validity of the Lagrangian information.
Overall, the EuLag Block can fully utilize the complementary advantages of Eulerian and Lagrangian perspectives in describing fluid dynamics, thereby being better suited for fluid prediction. For more implementation details of the EuLag Block, please refer to Appendix A.3.